if 2 meteroites in free space have masses of 10 kg and 15 kg respectively. find the of their center of mass if they are 12m apart.


1
Expert's answer
2021-02-10T21:37:17-0500
"CM=\dfrac{m_1r_1+m_2r_2}{m_1+m_2},""CM=\dfrac{10\ kg\cdot0+15\ kg\cdot12\ m}{10\ kg+12\ kg}=8.18\ m."

The center of mass is 8.18 meters from the 10 kg meteorite.